OpenStudy (anonymous):

9^2 + (c-1)^2 My book says its 81 + C^2 - 2c +1 How is that correct? Why wouldnt it just be 81^2 + C^2 +1 ?? It makes no sense to me? (c-2)^2 i thought was just C^2 +1 Please help

OpenStudy (anonymous):

Crap the bottom part i typed wrong i thought (c-1)^2 was C^2 +1

OpenStudy (a_clan):

(a-b)^2 = a^2 - 2*a*b - b^2

OpenStudy (anonymous):

When multiplying binomials use FOIL (First, Outside, Inside, Last) so for (c-1)^2 this is the same as (c-1)*(c-1) multiply the First terms of each binomial so c*c=c^2 multiply the Outside terms of each binomial so c*(-1)=-c multiply the Inside terms of each binomial so c*(-1)=-c multiply the Last terms of each binomial so (-1)*(-1)=1 Then add them all together to get: c^2+(-c)+(-c)+1 which simplifies to: c^2-2c+1 so 9^2 + (c-1)^2=81+c^2-2c+1

OpenStudy (anonymous):

How is that possible if (c-1)^2 is the same as c-1 * C-1 And the powers to powers rule says we multiply the eponets although there is no eponet so its assumed its 1

OpenStudy (anonymous):

ok so that is technically a binomial i understand that and foil. So your saying when you have multiple terms in pareth raised to a power you can not use the powers to powers rule and just multiple the exponets

OpenStudy (anonymous):

put in numbers and try it out: say c=3 (3-1)^2=2^2=4 if (c-1)^2=c^2+1 it would be 3^2+1=10, this is not correct however, with foil it is c^2-2c+1=3^2-2(3)+1=9-6+1=4, correct

OpenStudy (anonymous):

yep, you can only use the powers to powers rule with a single number, for a binomial you have to use FOIL

OpenStudy (anonymous):

Thank you!